OpenStudy (anonymous):

12 is 300% of? how do I work this problem?

OpenStudy (anonymous):

\[12=3.00 \times x\]

jhonyy9 (jhonyy9):

300 % --- 12 100 % --- x ---------- x=12(100)/300 =1200/300 = 4 - so hence 12 is 300% of 4